Instrument Log 

MPL-POL      C1       Lamont, OK Optical Trailer

Date

Operator

Comments

5/17/16

KT

Computer software patch updates were installed on the computer which caused the program to be restarted. At 1300 GMT the computer program was restarted. RTS at 1315 GMT.

12/31/2015

MS

Laser energy was not displayed nor cloud heights, and an alarm was beeping as well.  Power outage due to ice storm on 12/28/2015,  0255 GMT, some data will be missing.  Power was restored around 0200 GMT 12/31/2015.  Restarted program at 1501 GMT 12/31/2015.  RTS at 1526 GMT 12/31/2015.

11/17/2015

MS

A 7 hour power outage starting at 0700 GMT and power was returned at 1435 GMT, 11/17/2015. Instrument was down. During that time some data might be lost?  Started program back up and RTS at 1556 GMT, 11/17/2015.

9/28/15

CAM

The MPL laser power supply failed at approximately 1530 GMT on 9/24. Tried using the power supply from MPL102 on Friday, 9/25, but found the laser itself had also failed as there was no beam being transmitted.  On Monday, 9/28, swapped out MPL104 with MPL102.  Since the calibration and configuration files were on a physical PC (MPL104 had been on a virtual PC), the computer was installed as well.  RTS at 1923 GMT.  MPL104 will be returned to vendor for repair.

*Old System 104:  Transceiver: Model MPL-4B-IDS-532/POL-FS, SN 104.  Laser Power Supply: Model DC10H/532/D/R, SN 06-109D/R.  Data System: Model: MPL-4B-IDS-532/POL-FS, SN 104/4101.  Computer (was not actually in use): Dell Optiplex 380, SN 4CWDMN1, CSI 0343550. 

*New System 102:  Transceiver: Model MPL-4B-IDS-532/POL-FS, SN 102, WD41589.  Laser Power Supply: Model DC10H-532SS/D, SN 04-168/D, WD41590.  Computer: Dell Optiplex 990, SN B1ZPPS1, CSI 0346768. 

9/24/15

KT

The Laser power supply seemed to be off or not operating. There were no LED's lit, the laser beam was not visible during checks. At 1340 GMT power cycled the power supply and there was no changed in status. Contacted instrument mentor and notified him of what was found. Waiting to hear back from mentor on further troubleshooting. RTS at 1355 GMT.

7/28/15

KT

Instrument mentor reported that the MPL display was not accurate and requested that the instrument be inspected. During inspection it was found that the MPL window cover was still installed. Removed the cover and the data display returned to normal. RTS at 1915 GMT.

6/22/2015

MS

The graphs for the Sigma MPL and the Cross POL chart were not on at 0600 GMT 6/20/2015.  Reset both graphs for the MPL at 1319 GMT 6/22/2015, RTS at 1334 GMT.

6/10/15

CAM

Power to all ground mounted transformers is being cut in order to connect the new power panel near the AOS trailer and run power to new PGS building.  Instrument needs to be shut down.  Shut MPL down at 1349 GMT.  Restarted once power was restored and RTS at 1520 GMT.

6/4/15

CAM

Power to all ground mounted transformers is being cut in order to upgrade the transformers at AOS trailer, IDP2 and 50 profiler pad.  Instrument needs to be shut down.  Shut MPL down at 1306 GMT.  Restarted once power was restored and RTS at 1830

4/21/15

CAM

The computer needs to be rebooted following scheduled software and security patch updates.  Stopped program at 1828 GMT and rebooted computer. Restarted and RTS at 1831 GMT.

2/19/15

KT

Computer IT department uploaded patches and updates to the instrument computer. The computer needed to be rebooted. At 1520 GMT the computer was rebooted. At 1526 GMT the computer came online and was operational. RTS at 1535 GMT.

12/16/14

CAM

No data had been collected from MPL since 1811 GMT.  As per BCR-02040, all SGP Windows computers were updated and rebooted today.  This required a manual restart of MPL software.  Restarted SigmaMPL and RTS at 2154 GMT. 

9/26/14

KT

SDS reported no files had been collected from the instrument for 3 hours. At 1633 GMT found the computer program was locked up. Closed the program and was not able to get the program running. While working with Tim on the virtual computer we were able to get the instrument computer back online and operational. RTS at 1714 GMT.

9/24/14

CAM

All site power beyond the main office building needed to be shut down briefly to facilitate installation of an electrical panel at the old 50RWP pad.   IT department powered off the network switch at 1346 GMT in preparation for outage, which caused the instrument to lose connection.  Instrument was restarted and connection reestablished once power was restored.  RTS at 1714 GMT.

7/1/14

CAM

No data files had been collected since 1811 GMT.  Found SigmaMPL program was open but had stopped updating at 1841 GMT and showed error: Can’t Open AMCS Card: 2.  USB Server had lost connection with the data system.  Connected USB Sever to data system and SigmaMPL began updating and resumed normal operation at 1954 GMT.

6/24/14

CAM

The network went down at approximately 1800 GMT for planned upgrades.  This caused the USB server to lose connection with the data system and thus SigmaMPL program stopped.  Reconnected USB server to data system and restarted SigmaMPL software at 2101 GMT. 

6/9/14

CAM

No data files had been collected since 2211 GMT on 6/6.  Found SigmaMPL program was open but showed error: Can’t Open AMCS Card.  USB Server had lost connection with the data system.  Connected USB Sever to data system.  Restarted SigmaMPL software and RTS status O.K. at 1218 GMT.

6/2/14

6/3/14

CAM/TMG

MPL needs to be switched over to virtual computer.  Stopped program at 1808 GMT on 6/2.  Moved laser power supply serial connection to a Moxa device (data system was previously moved to a USB over IP server).  Installed SigmaMPL2013 software on virtual computer and pointed software to proper calibration files.  Restarted program and RTS at 1242 GMT on 6/3. 

5/6/14

CAM/TMG

Need to test USB over IP device server in preparation for running the MPL on a virtual computer.  Stopped program at 1453 GMT and connected data system to Startech USB over IP device server.  Computer was able to successfully communicate with data system.  Will leave in this configuration for a few days to test. RTS at 1508 GMT.

5/5/14

CAM

The benchtop surface of the bumpout where the MPL sits needs to be painted.  Shut MPL down at 1741 GMT and removed from the area so the surface could be painted.  Reinstalled MPL and RTS at 2106 GMT.

4/14/15

CAM/TMG

Need to test AnywhereUSB hub in preparation for running the MPL on a virtual computer.  Data acquisition was interrupted twice during testing, from 1838 to 1902 GMT and from 2019 to 2043 GMT.  RTS at 2043 GMT as previously configured. 

3/4/14

KT

System shut down due to power outage that occurred on 3/2/2014 at 2120 GMT. CF was closed 3/3/2014 due to snow storm. CF opened on 3/4/2014. Powered on the computer. System came online at 1420 GMT. RTS at 1435 GMT.

2/11/14

CAM

No data files had been collected since 1711 GMT.  Found SigmaMPL program was open but had locked up at 1626 GMT.  Closed and restarted SigmaMPL, but program would not start.  Rebooted computer and was then able to start SigmaMPL successfully.  RTS status O.K. at 2246 GMT.

2/10/14

CAM

No data files had been collected since 1711 GMT.  Found SigmaMPL program was open but had locked up at 1608 GMT.  Closed and restarted SigmaMPL, but program would not start.  Rebooted computer and was then able to start SigmaMPL successfully.  RTS status O.K. at 2131 GMT.

2/6/14

CAM

No data files had been collected since 1711 GMT on 2/5.  Found SigmaMPL was running, but the Save Data box was not checked under the Hardware tab.  Checked Save Data box and confirmed a data file was then created in ftphome.  RTS at 1342 GMT. 

2/4/14

CAM

No data files had been collected since 0111 GMT.  Found SigmaMPL was running, but the Save Data box was not checked under the Hardware tab.  Checked Save Data box and confirmed a data file was then created in ftphome.  RTS at 1354 GMT. 

1/22/14

CAM

No data files had been collected since 2311 GMT on 1/21.  Found SigmaMPL program was open but had locked up at 2232 GMT on 1/21.  Closed and restarted SigmaMPL, but program would not start.  Rebooted computer and was then able to start SigmaMPL successfully.  RTS status O.K. at 1311 GMT.

1/21/14

CAM

No data files had been collected since 1611 GMT on 1/17.  Found SigmaMPL program was open but had locked up at 1508 GMT on 1/17.  Closed and restarted SigmaMPL, but program would not start.  Rebooted computer and was then able to start SigmaMPL successfully.  RTS status O.K. at 1331 GMT.

1/7/14

CAM

No data files had been collected since 1511 GMT on 1/6.  Found SigmaMPL program was open but had locked up at 1459 GMT on 1/6.  Closed and restarted SigmaMPL but program would not start.  Rebooted computer and was then able to start SigmaMPL successfully.  RTS status O.K. at 1329 GMT.

11/27/13

KT

The program was not running. Due to patch updates it conflicted with the computer and would not allow the program to run. Contacted IT department and the issue was resolved. RTS at 1435 GMT.

10/4/13

JEM

The MPL has not been transferring any data. The MPL network cable got moved and the jack it was plugged into did not have access to the .98 network.  Connected that jack to the .98 network at the GIF switch.  RTS @1345 GMT.

10/2/13

CAM

MPL104 has returned from vendor and needs to be reinstalled.   Upgrades to 104 included incorporating the polarizer function with the transceiver and thus eliminating the external polarizer.  Stopped program at 1540 GMT and removed MPL 102 system (*transceiver, **laser power supply).  Installed new MPL *transceiver, **laser power supply and *** data system and connected to existing computer.  Renamed old SigmaMPL directory as SigmaMPL_10.2.13 and installed new software version SigmaMPL2013 to SigmaMPL directory.  Powered on and RTS at 1600 GMT, but not saving data at this time.  Notified mentor and he will copy the necessary calibration files before starting to save data. 

*Old - 102   *New – 104 

**Old - 04-168/D (WD41590)  **New - 06-109D/R  ***New – 104/4101 

9/17/13

CAM

Mentor requested more horizontal data be taken for MPL 102.  Stopped program at 2211 GMT.  Pointed MPL horizontally out the side window.  Collected horizontal data from 2229 to 0351 GMT.  Returned MPL to normal operating mode and RTS at 0354 GMT on 9/18.  Two horizontal data files (201309172229.mpl and 201309172300.mpl) were collected by SDS and will need to be flagged as invalid.  All other horizontal data files were moved out of ftphome before they were collected by SDS.

9/10/11

CAM

More evening horizontal data is needed for MPL 102.  Stopped MPL program at 2200 GMT and pointed MPL horizontally out the side window.  Noticed soon after that the farmer directly north of the site had started plowing the field, resulting in large dust cloud directly in the beam’s path.  Returned MPL to normal operating mode and RTS at 2232 GMT.  Will try again to get evening horizontal data when conditions are right.

9/4/13

CAM

Mentor requested more horizontal data be taken for MPL 102.  Stopped program at 0118 GMT.  Pointed MPL horizontally out the side window.  Collected horizontal data from 0127 to 0220 GMT.  Farmer had begun baling grass straight north of site, which was affecting the data.  Pointed MPL slightly more east and resumed collecting horizontal data from 0220 GMT to 0400 GMT.   Returned MPL to normal operating mode and RTS at 0405 GMT.  Moved all horizontal data out of ftphome directory so data will be missing between 0118 GMT and 0405 GMT.

8/2/13

CAM

Mentor requested horizontal data be taken for MPL 102.  Stopped program at 0041 GMT.  Pointed MPL horizontally out the side window.  Collected horizontal data from 0053 to 0201 GMT.  Returned MPL to normal operating mode and RTS at 0205 GMT.

7/22/13

CAM

Instrument mentor requested afterpulse and dark count data be collected for MPL 102.  Stopped program at 2040 GMT.  Covered telescope and collected afterpulse data from 2044 to 2059 GMT.  Shut laser off and collected dark count data from 2102 to 2118 GMT.  Restarted program in normal mode and RTS at 2121 GMT.  Moved requested files out of ftphome directory to C:\MPL\Cal_Data and renamed them appropriately.  SDS data will be missing data between 2040 and 2121 GMT.

7/9/13

CAM

Mentor requested that the MPL be replaced with the spare, which has just returned from the vendor.  Stopped program and removed MPL system (*transceiver, **laser power supply, *** data system, ****polarizer) at 1302 GMT.  Packed old MPL system (less the computer) and will send to vendor for repair.  Installed new MPL *transceiver and **laser power supply and connected to existing computer.  Renamed old SigmaMPL directory as SigmaMPL-SN104 and installed new software version SigmaMPL2013 to SigmaMPL directory.  Powered on and RTS at 1600 GMT.  *Old – 104   *New – 102  **Old - 06-109D/R  **New – 04-168/D (WD41590)  ***Old – 104/4101  ****Old - 104

5/15/13

CAM

There was no graphical data being displayed on the MPL computer.  ARM Diagnostic Plot Browser also stopped showing background signal for co pol and std co pol at around 1800 GMT on 5/13.   Stopped SigmaMPL program at 1915 GMT and rebooted computer.  Restarted SigmaMPL program and RTS status O.K. at 1917 GMT.  Data is now being plotted as it should.

4/15/13

CAM

No data files had been collected since 2111 GMT on 4/14.  Found MPL software had scrolling error: Card 0 Timeout.  Stopped software and power cycled electronics and rebooted computer.  RTS status O.K. at 1222 GMT.

4/8/13

CAM

The MPL corrector plate needed to be cleaned.  Stopped data acquisition at 1826 GMT.  Powered laser off and tilted telescope horizontal and cleaned.  Returned telescope to normal position and restarted data acquisition at 1858 GMT.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Powered UPS off at 1730 GMT.  Power was restored at 1920 GMT.  Powered UPS on at 1945 GMT.  Instrument and computer were restarted at 2000 GMT on 2/26 but laser shutter was not opened.  Opened laser shutter at 2133 GMT on 2/27.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  MPL was gracefully shut down at 1351 GMT in anticipation of power outage.  Power was restored at 2115 GMT after electrical upgrades were complete.  Restarted MPL and RTS at 2142 GMT.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  Restarted MPL and RTS at 2052 GMT.

10/18/12

CAM

Computer hard drive failed on 10/16 and was replaced by IT department on 10/17.  Computer needs to be reinstalled.  Reinstalled *computer at 1700 GMT.  MPL software would not load drivers and run properly.  Notified mentor of problem.  Mentor replaced AMCSUSB.dll in C:\MPL and added Cygwin programs for MPL support in /usr/local/bin.  RTS at 2000 GMT. 

*New – 4CWDMN1  Model - Optiplex 380  CSI# 0343550

10/17/12

CAM

No data had been collected since 2211 GMT on 10/16.  Found computer at blue screen with kernel_stack_inpage_error displayed.  Power cycled computer but it would not boot up.  Attempted to boot to safe mode which also failed.  Computer freezes at initial splash screen and becomes unresponsive.  Disconnected power for several hours and tried again with the same results.  Removed *computer and took to computer lab for repair or replacement.  Instrument remains down at this time.  *Old – 4CWDMN1  Model - Optiplex 380  CSI# 0343550

7/31/12

KT

The MPL window needed cleaning. At 1802 GMT stopped the program and powered down system. The MPL window was cleaned. At 1850 GMT powered on system.

11/7/11

CAM/JEM

Received replacement telescope temperature sensor and need to install.  Removed faulty telescope temperature sensor at 1804 GMT.  Removed heat shrink from cable/sensor and found one lead of the sensor was broken. Removed faulty sensor and soldered new sensor to cable.  Stopped MPL program at 1917 GMT and tipped the telescope sideways in order to epoxy new sensor to the telescope.  Left telescope in this position overnight to allow epoxy to dry.  Returned telescope to normal position and restarted program the following morning.  RTS at 1422 GMT on 11/8/11.

9/28/11

CAM

DQPR 2723 indicates: Scope temp intermittently jumps up over 4x10^7 ºC between 5/3 and 5/5, but then is consistently at that level on 5/13 through current.  T/S found temperature sensor (thermocouple) is faulty where it is epoxied to telescope. Taped sensor cable securely to telescope at 1437 GMT and reading returned to normal.  Mentor will order replacement sensor.

8/10/11

PAD

Computer was replaced on 8/9/11 but replacement computer is not active.  Computer department removed computer and moved hard drive from failed computer into new one.  Installed new computer at 1641 GMT.  Computer and instrument started correctly and is now operational.  Old Model DCCY, SN FG 72281.  New Model Optiplex 380, SN 4CWDMN1, CSI 0343550.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.  Power was restored on 8/9/11 at 0211 GMT.  At 1800 GMT, computer will not boot up.  Removed computer and took to computer lab.  TS found computer power supply had failed.  Computer department provided replacement computer.  Installed computer with computer department and booted up.  Program started correctly.  RTS at 1927 GMT.

4/29/11

JEM

The MPL window is loose and could leak.  The cover was installed from 1624 to 1628 while the inside and outside edges of the window was sealed.

4/25/11

MK

Reinstalling the MPL that has returned from vendor repair.  Installed Telescope S/N 104, Lidar data system Model - MPL-4B-532/Pol-FS S/N -104/4101, Laser controller/power unit DC10H/532D/R S/N 06-109D/R, Pol control interface DR951483, Dell computer DCC4 S/N FG72281. Powered on instrument at 1530 GMT, contacted Instrument mentor.

2/4/11

MK

Due to blizzard-like conditions we had at the site this week, fine powdery snow has come in through the MPL bumpout thus causing serious water intrusion on top of the MPL corrector plate. The water has made its way inside the telescope. The MPL was powered off and taken partially apart so the electronics could dry out.  The MPL will need to be shipped back to Sigma Space for cleaning, inspection and testing.  Disconnected and removed all MPL components. Telescope S/N 104, Lidar data system  Model - MPL-4B-532/Pol-FS  S/N -104/4101, Laser controller/power unit  DC10H/532D/R  S/N 06-109D/R, Pol control interface DR951483, Dell computer DCC4 S/N FG72281.  2/4/11

1/25/11

CAM/PAD

Corrector plate has collected dirt/dust and unable to clean sufficiently with canned air.  Consulted with mentor and was advised to clean corrector plate.  Stopped program and shut down laser.  Cleaned corrector plate as per mentor's instructions.  RTS at 2258 GMT.

12/1/10

PAD

AC power to CF will be shut down today for ongoing construction.  Instrument needs to be stopped until power is restored.  Shut down instrument at 1433 GMT.  Restarted at 1837 GMT.

11/18/10

CAM

AC power needed to be secured to drill pier holes for new construction near the RCF trailer.  Need to gracefully shut the instrument down prior to power outage.  Powered instrument down at 1441 GMT.  Restarted instrument after AC power was restored.  RTS at 1803 GMT.

11/10/10

CAM

AC power will be interrupted for construction near the main compound.  Need to gracefully shut the instrument down prior to power outage.  Powered instrument down at 2134 GMT.  Restarted instrument after AC power was restored.  RTS at 2305 GMT.

7/19/10

CAM

Instrument mentor requested the telescope temperature sensor be moved from channel 1 to channel 2 of the A/D converter.  Stopped the program at 2006 GMT and powered the system off.  Removed the transceiver cover and moved the sensor from channel 1 to channel 2.  Replaced cover, restored power and RTS at 2034 GMT.

7/9/10

PAD

Last SDS collection was at 0911 GMT.  MPLPOL computer has locked up several times over the last week.  Computer needs to be replaced.  Removed *computer at 1455 and took to computer department so harddrive could be swapped.  Installed new computer.  RTS at 1525 GMT.  *Old Dell GX620 USFF Computer: SN# 28WX181 CSI 0332123.  *New Dell GX620 SFF Computer: SN# FG72281 Los Alamos 1153893.

7/7/10

CAM

No data had been collected from the MPL since 2111 GMT on 7/6.  Found the computer had encountered a problem and Windows had shut down.  Restarted computer and program RTS at 1238 GMT.

7/6/10

PAD

Computer department was remotely updating windows updates and are now unable to remotely connect with MPL computer.   Computer has power but nothing is shown on monitor.  Rebooted computer and program restarted correctly.  Computer department was able to connect and update software.  RTS at 1752 GMT.

7/6/10

PAD

Last SDS collection was on 7/3/10 at 1711 GMT.  Computer has power but nothing is shown on monitor.  Rebooted computer and program restarted correctly.  RTS at 1344 GMT.

6/24/10

PAD

Computer has power but is not showing anything on monitor.  Powercycled computer and program restarted correctly.  RTS at 1451 GMT.

6/9/10

PAD

Last SDS collection was on 6/8/10 at 2211 GMT.  Computer has power but nothing is shown on monitor.  Power cycled computer and program started correctly.  RTS at 1303 GMT.

6/7/10

PAD

Last SDS collection was at 0411 GMT.  Computer has power but nothing is shown on monitor.  Power cycled computer and program started correctly.  RTS at 1555 GMT.

6/2/10

CAM

Computer department needs to install software updates on the MPL-POL-FS computer.  Shut the MPL down at 1446 GMT and removed the computer to the computer lab for updating.  Reinstalled the computer and MPL-POL-FS RTS at 1843 GMT.

6/1/10

CAM

Need to install the new upgraded fast switching *MPL-POL (MPL-POL-FS)  The old MPL-POL will then be sent in to be upgraded as well. Installed the MPL-POL-FS next to the old system and powered on at 1615 GMT.  The old MPL-POL was left running at this time.  Computer department later added Service Pack 3 to the new computer and set the IP address to be that of the old.  Moved network cable from the old MPL computer to the new at 2003 GMT.  Powered the old MPL-POL off at 2004 GMT.  Will pack up the old MPL and ship to Sigma Space. 

*Old – 106  *New - 104

2/16/10

TMG

MPL has not been transferring files since the new computer was installed.  Data files are being created, just not transferred.  Stopped MPL program at 2057 GMT.  Shutoff sigwin service and installed Filezilla FTP server and setup the SDS account.  Rebooted computer and RTS at 2121 GMT.

2/16/10

CAM

No data files had been received from the MPL since 1711 GMT on 2/14.   Computer power LED was lit yellow but nothing was displayed on the monitor.  Computer was unresponsive to mouse or keyboard activity.  Power cycled the computer several times with no change in status.  Computer would not power up.  Removed *computer and took to computer lab for diagnostics.  Suspect failed motherboard.  Removed the hard drive and installed in a new computer.  Installed new computer and reloaded USB component drivers.  Restarted program and RTS status O.K. at 1829 GMT.

*Old - - 2R1QJ71 (Dell Optiplex GX280, WD 54903) 

*New - 1F72281 (Dell Optiplex GX620, US Gov 1153895 Los Alamos)

2/3/10

CAM

SDS instrument status page indicated no files had been received from the MPL since 2011 GMT.   Computer power LED was lit but nothing was displayed on the monitor.  Computer was unresponsive to mouse or keyboard activity.  Power cycled the computer at 2156 GMT.  Program restarted and RTS status O.K. at 2158 GMT. 

1/15/10

JEM

Replaced failed room air circulation fan at 2215 GMT.

1/12/10

CAM

The small fan that is used to help circulate room air around the MPL in the bumpout area has failed.  Will place an order for a new fan.

1/12/10

CAM

SDS instrument status page indicated no files had been received from the MPL since 1811 GMT.   Computer power LED was lit but nothing was displayed on the monitor.  Computer was unresponsive to mouse or keyboard activity.  Power cycled the computer at 2052 GMT.  Program restarted and RTS status O.K. at 2054 GMT. 

11/23/09

PAD

Last SDS collection was on 11/21/09 at 1611 GMT.  Computer has power but is not active.  Nothing is shown on monitor.  Beam is being emitted by instrument.  Power cycled computer and verified computer and program started correctly.  RTS at 1534 GMT.

11/10/09

PAD

Last SDS collection was at 1211 GMT.  Found power strip switch off that supplies power to GIF trailer network hub.  Powered up hub and communications were established.  Able to ping SDS collector from computer but SDS is still unable to collect data.  Verified data was on computer.  Rebooted computer at 2005 GMT and computer department was able to force data collection.  RTS at 2045 GMT.

10/19/09

PAD

Contractors on site to apply latex sealant to GIF building roof.  Placed cover on instrument and covered MPL bay window at 1759 GMT.  Removed cover at 2007 GMT. 

7/1/09

PAD

Computer has power but is not active or responsive.  Power cycled computer and verified program restarted correctly.  RTS at 1251 GMT.

6/25/09

TMG

Computer software needs updated.  Stopped program @ 1552 GMT.  Installed Service Pack 3 @ 1555 GMT.   Rebooted and RTS @ 1635 GMT.

1/27/09

CAM

Instrument mentor advised that the background level has been increasing for some time now and requested the detector be replaced.  Placed the cover over the telescope at 1837 GMT to obtain pre-replacement dark counts.  Ran with the laser on from 1837 to 1842 GMT.  Turned laser shutter off and ran with telescope covered and laser off from 1842 to 1847 GMT.  Powered system off at 1847 GMT.  R/R *detector.  Powered on and ran system with telescope covered and laser off from 1920 to 1925 GMT to obtain post-replacement dark counts.  Powered laser on and ran with telescope covered until 1930 GMT.  Uncovered telescope, rebooted computer and RTS in normal operational mode at 1932 GMT.   *Old - 13970-1 (model # SPCM-AQR-14-FC)    *New – 16494 (model # SPCM-AQRH-14-FC)

4/17/08

Mentor

Upgraded SGP MPL 106 from SigmaMPL V2.20 to SigmaMPL V2.27 with autostart and laser manage.  The system was stopped/down for about 1 minute at 2010 GMT.

2/5/08

CAM

Noticed the fan that has been used for circulating room air into the bumpout was now sitting on the floor and not powered on.  Notified instrument mentor who requested the fan be placed back into service. Set fan back in bumpout and powered on at 2020 GMT.  It’s not clear how long the fan has been off.

12/11/07

Mentor

Bin resolution changed from 30m to 15m at 2100 GMT as per BCR 01438.

11/13/07

PAD

Last SDS collection was on 11/9/07 at 2210 GMT.  Computer and instrument has power.  MPL program is not running.   Program was stopped on 11/9/07 and new time server software was downloaded.  Program was restarted at that time but did not start correctly.  Restarted program and power cycled digital interface box.   RTS at 1508 GMT.

11/5/07

CAM

Instrument mentor contacted Site Ops and indicated the MPL computer had locked up around 2045 GMT while a driver was being updated remotely.  Checked system and found the computer was at the Windows blue screen memory dump error.  Rebooted computer and contacted mentor who remotely reloaded the old driver.  Restarted MPL program and RTS at 2103 GMT.

10/19/07

CAM

Instrument mentor requested that the polarizer be power cycled in order to determine what version of software it is running.  Power cycled the polarizer a couple of times at 2100 GMT and determined from the light sequence that it is running V2.05..  After the power cycle, the polarizer was not alternating between modes.  Tried stopping and restarting the MPL program but the polarizer still was not alternating.  Finally rebooted the computer and restarted the program and then polarizer was working properly.  RTS at 2106 GMT.

10/10/07

CAM

The MPL has been reporting excessively high dark counts.  It was previously determined that the detector had failed.  A new detector was supplied by Sigma Space to replace the failed one under warranty.  Placed the cover over the telescope at 1433 GMT to obtain pre-replacement dark counts.  Ran with the laser on from 1433 to 1438 GMT.  Turned laser shutter off and ran with telescope covered and laser off from 1438 to 1443 GMT.  Powered system off at 1444 GMT.  R/R *detector.  Also switched the polarizer over to USB port and connected the laser power supply to COM1 as requested by mentor.  Powered the computer on at 1520 GMT.  Mentor made necessary software adjustments to account for cabling changes.  Ran system with telescope covered and laser off from 1552 to 1555 GMT to obtain post-replacement dark counts.  Powered laser on and ran with telescope covered until 1600 GMT.  Uncovered telescope and RTS in normal operational mode at 1602 GMT.   *Old - 12862     *New - 13970-1

10/3/07

CAM

Site power was lost for an extended time yesterday and system had to be shut down.  Power was restored last night after normal business hours.  System needs to be restarted.  Restarted the MPL at 1339 GMT.  RTS status O.K.

10/2/07

CAM

The site lost power around 1940 GMT during a severe thunderstorm. Kay Electric advised it could be off for some time due to a failed transformer near Billings.  Performed a graceful shutdown of the MPL at 2010 GMT.  Will restart when site power is restored.

9/18/07

PAD

Computer was restarted on 9/14/07 at 1458 GMT after it had last SDS collection on 9/13/07.  Digital interface box was not power cycled at that time and indicator light is not blinking.  Power cycled digital interface box at 1522 GMT.  RTS at 1537 GMT.

9/14/07

PAD

Last successful SDS collection was on 9/13/07 at 2310 GMT.  Computer has power but nothing is shown on monitor.  Computer is not responsive to keyboard or mouse.  Power cycled computer at 1458 GMT and restart MPL program.  RTS at 1513 GMT.

8/24/07

PAD

Last successful SDS collection was at 0210 GMT.  Computer and instrument has power.  Laser is emitting beam but computer is not responsive to keyboard or mouse.  UPS that supplies power to instrument and computer is in alarm.  We are experiencing power bumps at this time.  Power cycled computer and digital interface box.  Notified facilities of UPS alarm.  RTS at 1411 GMT.

8/6/07

CAM

The telescope temperature is reading over 60C.  Checked the sensor cable and connector and found the SMA connector was loose.  Movement of the cable caused the temperature reading to vary tremendously.  Tightened the SMA connector and temperature reading dropped to around 23C and was stable.  Also returned the fan to the bumpout so that it is circulating air around the instrument.  RTS at 1927 GMT.

8/3/07

MK

The MPL computer was found locked up.  Power cycled computer and digital interface box at 1508 GMT.  System returned to normal.

7/13/07

PAD

Last SDS collection was at 0410 GMT.  Found computer and all components have power.  Laser is emitting.  Computer is not responsive to keyboard or mouse.  Digital interface status box is not flashing.  Power cycled computer and digital interface box.  Restarted program.  Notified instrument mentor.  RTS at 1450 GMT.

7/3/07

MK

The scope_temp is running too high.Checked cable connections and power cycled the polarization interface. Temperature dropped to approximately 60. RTS at 1950 GMT.

6/14/07

PAD

Last SDS collection was at 0210 GMT.  Computer has power but is not responsive to keyboard or mouse.  Power cycled computer and digital interface box.  Restarted MPL program.   RTS at 1423 GMT.

4/27/07

CAM

The polarizer controller was locked up and not alternating.  Power cycled the polarizer controller at 1708 GMT.  RTS status O.K.

4/24/07

PAD

Last SDS collection was at 0110 GMT.  Computer has power but is not responsive to keyboard or mouse.  Power cycled computer at 1357 GMT.  Restarted MPL program.  RTS at 1412 GMT.

4/3/07

MK

The MPL is locked up with black screen. Powered off computer and restarted at 1510 GMT. Also power cycled polarization module. RTS at 1530 GMT.

3/13/07

PAD

Last SDS collection was at 0210 GMT.  Found computer has power but is not responsive to keyboard or mouse.  Power cycled computer and restarted program.  RTS at 1437 GMT.

2/20/07

PAD

Computer has power but nothing is shown on monitor and is not responsive to mouse or keyboard.  Power cycled computer and restarted program.  RTS at 1540 GMT.

1/8/07

PAD

Last successful collection was at 0010 GMT.  Computer has power but nothing is shown on monitor and computer is not responsive.  Power cycled computer at 1516 GMT.  Restarted program.  RTS at 1531 GMT.

12/18/06

CAM

SDS instrument status page indicated no files had been received from the MPL since 1710 GMT on 12/15.  Collection log says ‘Failed to get new FTP object’.  Checked instrument and found it was running and generating data files in ftphome directory as it should.  Rebooting the computer did not resolve collection problem.  Notified Ron Reed, who found that someone had initiated a file transfer on 12/15 and that appeared to have hung the data collections.  Corrected the problem, restarted the software and RTS at 1926 GMT.

12/18/06

PAD

Last successful collection was on 12/15/06 at 1710 GMT.  Files are on computer and communicating with SDS.  Rebooted computer and restarted program at 1532 GMT.  This did not fix collection problem.  Contacted computer department.  Troubleshooting at this time.

12/12/06

PAD

Last SDS collection was on 12/11/06 at 2210 GMT.  Computer has power but nothing is shown on monitor.  Computer is not responsive.  Power cycled computer and digital interface box.  Restarted Sigmampl program.  Program appears to have restarted correctly.  RTS at 1508 GMT.

11/21/06

PAD

Last successful collection was on 11/20/06 at 2310 GMT.  Computer has power but is not responsive.  Laser is on and is transmitting beam.  Power cycled computer and digital interface box.  Restarted MPL program.  Program restarted correctly.  RTS at 1518 GMT.

10/31/06

PAD

Last successful collection was at 0210 GMT.  Computer had power but was not responsive.  Power cycled computer and digital interface at 1544 GMT.  Restarted MPL program.  RTS at 1559 GMT.

10/10/06

PAD

Computer has power but monitor shows nothing.   Laser has power is transmitting a beam.  Rebooted computer at 1845 GMT.  Restarted MPL program.  RTS at 1900 GMT.

9/19/06

PAD

Last SDS collection was on 9/18/06 at 2210 GMT.  At 1414 GMT, found computer had power but was not responsive to keyboard or mouse.  Nothing is shown on monitor.  Power cycled computer at 1415 GMT and computer restarted correctly.  After starting MPL program, the scrolling status displays:  Polarizer not open.  I power cycled the interface box with same results.  At 1429 GMT, Chris Martin shut down computer, interface box, and laser power supply and restarted with same results.  Chris reseated all connectors on computer and instrument and restarted.  Program restarted correctly.  RTS at 1444 GMT.

8/29/06

CAM

MPL program is not running.  Started Sigma MPL 2.12 program at 1815 GMT.  RTS status O.K.  Suspect the computer had rebooted due to problems with the UPS.  A new UPS battery has been ordered.

8/29/06

PAD

UPS in GIF trailer is failing self diagnostics test.  Instruments are losing power.  Computer is running but MPL program is not.  Restarted MPL program at 1522 GMT.  RTS at 1537 GMT.  Program was started again at 1626 GMT. 

8/28/06

PAD

Storms in area had tripped UPS located in GIF trailer.  Dan Nelson reset UPS.  Restarted MPL program at 1418 GMT.  RTS at 1433 GMT.

7/31/06

MK

MPL computer appears to be locked up. Rebooted computer at 1445 GMT. Power cycled the polarization interface at 1520 GMT. Range is set to 30M.

8/9/06

CAM

Installed the new *Polarized MPL.  Placed into service and operational at 1614 GMT.  *New - 106